# Rethinking LLM Memorization through the Lens of Adversarial Compression

### v2

@@ 2 Do We Really Need Another Notion of Memorization?
    Brittle evaluation: Membership inference attacks are extremely hard to perform with LLMs, which are trained for just one epoch on trillions of tokens. Some recent work shows that LLM membership inference is extremely brittle (Duan et al., 2024; Maini et al., 2021).
@@ 3.1 A New Definition of Memorization
The threshold $\tau(y)$ is a configurable parameter of this definition. We might choose to compare the ACR to the compression ratio of the text when run through a general-purpose compression program (explicitly assumed not to have memorized any such text) such as GZIP (Gailly and Adler, 1992) or SMAZ (Sanfilippo, 2006). This amounts to setting $\tau(y)$ equal to the SMAZ compression ratio of $y$, for example. Alternatively, one might even use the compression ratio of the arithmetic encoding under another LLM as a comparison point, for example if it was known with certainty that the LLM was never trained on the target output, and hence could not have memorized it (Delétang et al., 2023). In reality, copyright attribution cases are always subjective, and the goal of this work is not to argue for the right threshold function, rather to advocate for the adversarial compression framework for arbitrating fair data use. Thus, we use $\tau=1$ in the experiments below, which we believe has substantial practical value. [^1] There exist prompts like “count from $1$ to $1000$,” for which a chat model $M$ is able to generate “$1,2,\ldots,1000$,” which results in a very high ACR. However, for copyright purposes, we argue that this category of algorithmic prompts are in the gray area where determining memorization is difficult and beyond the scope of this paper given our primary application to creative works.
@@ 3.2 MiniPrompt: A Practical Algorithm for Compressible Memorization
In all of our experiments below, when we present memorization metrics using compression, we are showing the results of running our MiniPrompt algorithm. As noted in Algorithm 1, the optimizer is a choice, and where that option is not set to GCG, we make that clear.
@@ 4.2 TOFU: Unlearning and Memorization with Author Profiles
We compare completion and compression tests on the TOFU dataset (Maini et al., 2024). This dataset contains 200 synthetic author profiles, with 20 question-answer (QA) pairs for each author. We finetune Phi-1.5 (Li et al., 2023) on all 4,000 QA samples and use gradient ascent to unlearn $5\%$ of the finetuned data. Following the TOFU framework (Maini et al., 2024), we finetune with a learning rate of $2\times 10^{-5}$ and reduce the learning rate during unlearning to $1\times 10^{-5}$. Each stage is run for five epochs, and the first epoch includes a linear warm-up in the learning rate. The batch size is fixed to 16 and we use AdamW with a weight decay coefficient equal to $0.01$.
@@ 4.5 Validation of MiniPrompt with Four Categories of Data
To further determine the validity of our definition, we investigate the average compression rate of natural text that is not in the training set. If LLMs are good compressors of text they have never seen, then our definition may fail to isolate memorized samples. We take random sentences from Associated Press articles that were published in November 2023, well after the models we experiment with were trained. These strings are samples from the distribution of training data as the training set includes real news articles from just a few months prior. Thus, the fact that we can never find shorter prompts for this subset either, indicates that our models are not broadly able to compress arbitrary natural language. Again, see the zero-height bar in Figure 6.
@@
Finally, we look at the memorization of training samples that are not common or famous, but that do exist in the training set. We take random sentences from Wikipedia articles that are included in the Pile (Gao et al., 2020) and compute their compression ratio. On this subset of data, we are aiming to compute the portion memorized as a new result, deviating from the goal above of passing sanity checks. Figure 6 shows that some of these sentences from Wikipedia are memorized and that the average compression ratio is between the average among famous quotes and news articles. Note that the memorized samples form this subset are strings that appear many times on the internet like ‘‘The Burton is a historic apartment building located at Indianapolis, Indiana.’’
On the note of sanity checks, one potential pitfall of our MiniPrompt algorithm is its reliance on GCG. It is possible that there exist shorter strings than we can find. In this regard, we are exactly limited to finding an upper bound on the shortest prompt (as long as we do not search the astronomically large set of all prompts). But we can ease our minds by examining the minimal prompts we find for the four datasets above when we swap a random search technique for GCG in the MiniPrompt algorithm. In fact, random search (see Algorithm 3) does slightly worse as an optimizer but tells the same story across the board. Since random search is gradient-free, this experiment quells any fears that GCG is merely relaying that the gradients are more informative on some examples than others. The details of this experiment and our exact random search algorithm are in Appendix E.
@@ Acknowledgments
Prompt Optimization
We borrow prompt optimization tools from work on jailbreaking where the goal is to force LLMs to break their alignment and produce nefarious and toxic output by way of optimizing prompts [Zou et al., 2023, Zhu et al., 2023, Chao et al., 2023, Andriushchenko, 2023]. Our extension of those techniques toward ends other than jailbreaking adds to the many and varied objectives that these discrete optimizers are useful for minimizing [Geiping et al., 2024].
@@
Рис. 7: Left: Fraction of Harry Potter texts that are compressible. Right: an example of hard tokens that elicit Harry Potter text.
@@ E.1 More Models and Discrete Optimizers
Рис. 8: Pythia-410M Memorization with GCG.
Рис. 9: Pythia-1.4B Memorization with Random Search.
@@ E.2 Alternative Thresholds
In the main body of this paper we discuss various choices for the threshold function $\tau$ and we continue that discussion here. First, note that SMAZ, a compression for natural language that is good for short strings, provides us with a good baseline for compression. In Figure 10, we show the Pythia-1.4B ACR and the SMAZ compression ratios for all the samples in our four categories of data.
Рис. 10: Comparing SMAZ compression ratios to the ACR according to Pythia-1.4B of four categories of data.

### v3

@@ 2 Do We Really Need Another Notion of Memorization?
In other words, this completion-based test is too conservative to capture memorization when model owners may take steps to make it look like their model has not memorized certain data. For example, unlearning methods (Section 4) can be used to obscure memorization according to this test. Our definition, which relies on optimization and not completion alone, is not fooled by these minor tricks to appear compliant.
@@
    Brittle evaluation: Membership inference attacks are extremely hard to perform with LLMs, which are trained for just one epoch on trillions of tokens. Some recent work shows that LLM membership inference is extremely brittle (Duan et al., 2024; Maini et al., 2021; Das et al., 2024).
@@ 3.1 A New Definition of Memorization
The threshold $\tau(y)$ is a configurable parameter of this definition. We might choose to compare the ACR to the compression ratio of the text when run through a general-purpose compression program (explicitly assumed not to have memorized any such text) such as GZIP (Gailly and Adler, 1992) or SMAZ (Sanfilippo, 2006). This amounts to setting $\tau(y)$ equal to the SMAZ compression ratio of $y$, for example. Alternatively, one might even use the compression ratio of the arithmetic encoding under another LLM as a comparison point, for example if it was known with certainty that the LLM was never trained on the target output, and hence could not have memorized it (Delétang et al., 2023). In reality, copyright attribution cases are always subjective, and the goal of this work is not to argue for the right threshold function, rather to advocate for the adversarial compression framework for arbitrating fair data use. Thus, we use $\tau=1$ in the experiments below, which we believe has substantial practical value. [^1] There exist prompts like “count from $1$ to $1000$,” for which a chat model $M$ is able to generate “$1,2,\ldots,1000$,” which results in a very high ACR. However, for copyright purposes, we argue that this category of algorithmic prompts are in the gray area where determining memorization is difficult and beyond the scope of this paper given our primary application to creative works. For more discussion on alternative thresholds, see Appendix E where we discuss the implications of this choice.
@@ 3.2 MiniPrompt: A Practical Algorithm for Compressible Memorization
In all of our experiments below, when we present memorization metrics using compression, we are showing the results of running our MiniPrompt algorithm. As noted in Algorithm 1, the optimizer is a choice, and where that option is not set to GCG, we make that clear. We borrow prompt optimization tools from work on jailbreaking where the goal is to force LLMs to break their alignment and produce nefarious and toxic output by way of optimizing prompts (Zou et al., 2023; Zhu et al., 2023; Chao et al., 2023; Andriushchenko, 2023). Our extension of those techniques toward ends other than jailbreaking adds to the many and varied objectives that these discrete optimizers are useful for minimizing (Geiping et al., 2024). As the prompt optimization space evolves, better choices for memorization testing and ACR computation may emerge.
@@ 4.2 TOFU: Unlearning and Memorization with Author Profiles
We compare completion and compression tests on the TOFU dataset [^2] This dataset is released under the MIT License and their assets and license can be found at https://huggingface.co/datasets/locuslab/TOFU. (Maini et al., 2024). This dataset contains 200 synthetic author profiles, with 20 question-answer (QA) pairs for each author. We finetune Phi-1.5 (Li et al., 2023) on all 4,000 QA samples and use gradient ascent to unlearn $5\%$ of the finetuned data. Following the TOFU framework (Maini et al., 2024), we finetune with a learning rate of $2\times 10^{-5}$ and reduce the learning rate during unlearning to $1\times 10^{-5}$. Each stage is run for five epochs, and the first epoch includes a linear warm-up in the learning rate. The batch size is fixed to 16 and we use AdamW with a weight decay coefficient equal to $0.01$.
@@ 4.5 Validation of MiniPrompt with Four Categories of Data
To further determine the validity of our definition, we investigate the average compression rate of natural text that is not in the training set. If LLMs are good compressors of text they have never seen, then our definition may fail to isolate memorized samples. We take random sentences from Associated Press [^3] We use data from the Associated Press and their terms of service allow for this use but not redistribution. Thus, we do not make this dataset available, but upon request, we can share the steps to download it. See their terms here: https://apnews.com/termsofservice. articles that were published in November 2023, well after the models we experiment with were trained. These strings are samples from the distribution of training data as the training set includes real news articles from just a few months prior. Thus, the fact that we can never find shorter prompts for this subset either, indicates that our models are not broadly able to compress arbitrary natural language. Again, see the zero-height bar in Figure 6.
One important thing to consider is that optimized prompts can elicit any output from the model (in particular things that were never seen during training). Our experiments, including those involving random strings and Associated Press data, show that while the GCG algorithm can elicit any output, we never observe compression of non-training data. This suggests that our method is robust against false positives.
@@
Finally, we look at the memorization of training samples that are not common or famous, but that do exist in the training set. We take random sentences from Wikipedia articles that are included in the Pile [^4] This dataset is released under the MIT License. (Gao et al., 2020) and compute their compression ratio. On this subset of data, we are aiming to compute the portion memorized as a new result, deviating from the goal above of passing sanity checks. Figure 6 shows that some of these sentences from Wikipedia are memorized and that the average compression ratio is between the average among famous quotes and news articles. Note that the memorized samples form this subset are strings that appear many times on the internet like ‘‘The Burton is a historic apartment building located at Indianapolis, Indiana.’’
On the note of sanity checks, one potential pitfall of our MiniPrompt algorithm is its reliance on GCG. It is possible that there exist shorter strings than we can find. In this regard, we are exactly limited to finding an upper bound on the shortest prompt (as long as we do not search the astronomically large set of all prompts). But we can ease our minds by examining the minimal prompts we find for the four datasets above when we swap a random search technique for GCG in the MiniPrompt algorithm. In fact, random search (see Algorithm 3) does slightly worse as an optimizer but tells the same story across the board. In other words, one might fear that our findings are the results of some peculiarity in GCG or some bias/preference GCG has for finding short prompts on some types of data. We establish that the same general trends in memorization can be observed with a gradient-free search algorithm, and thus conclude that we are not mistaking a GCG bias for some other signal. Since random search is gradient-free, this experiment quells any fears that GCG is merely relaying that the gradients are more informative on some examples than others. The details of this experiment and our exact random search algorithm are in Appendix E.
Рис. 7: ACR versus target string length. Our experiments are designed to include a balanced mix of both short and long sequences, ensuring that the evaluation of the ACR metric is comprehensive and unbiased. This helps mitigate any potential bias towards longer sequences.
Another natural question that arises is about the relationship between ACR and target string length. To address this issue, we plot the sequence length on the horizontal axis to illustrate how the ACR varies across different lengths in Figure 7. This analysis shows that while longer sequences can achieve higher compression ratios, the ACR metric remains effective and meaningful for shorter sequences as well.
@@ Acknowledgments
  • Das et al. (2024) Debeshee Das, Jie Zhang, and Florian Tramèr. Blind baselines beat membership inference attacks for foundation models. arXiv preprint arXiv:2406.16201, 2024.
@@
In our experiments we use GCG [Zou et al., 2023] so we provide a pseudoscope description of it here along with our MiniPrompt algorithm.
@@
Рис. 8: Left: Fraction of Harry Potter texts that are compressible. Right: an example of hard tokens that elicit Harry Potter text.
@@ E.1 More Models and Discrete Optimizers
Рис. 9: Pythia-410M Memorization with GCG.
Рис. 10: Pythia-1.4B Memorization with Random Search.
@@ E.2 Alternative Thresholds
In the main body of this paper we discuss various choices for the threshold function $\tau$ and we continue that discussion here. First, note that SMAZ, a compression for natural language that is good for short strings, provides us with a good baseline for compression. In Figure 11, we show the Pythia-1.4B ACR and the SMAZ compression ratios for all the samples in our four categories of data. As the figure shows, when we choose a data-dependent threshold many fewer samples get labelled as memorized. This is a reasonable knob for regulators and litigators to turn. In a court of law (in the United States), this kind of evidence would be more or less compelling, but in every case would still contribute to the evidentiary body presented and help lawyers make cases about copyright infringement.
Рис. 11: Comparing SMAZ compression ratios to the ACR according to Pythia-1.4B of four categories of data.
E.3 Paraphrased Famous Quotes
We also explore how compressible paraphrased versions of famous quotes are. This line of questioning aims to disentangle exact match memorization from concept memorization. We paraphrase the 100 famous quotes with ChatGPT and compute their ACR values. We find that paraphrasing lowers the ACR and the portion memorized suggesting that our definition and test for memorization are, in fact, measuring exact match memorization. See Table 1.
Таблица 1: Memorization statistics for Pythia-1.4B model.
Model Data Avg. ACR Portion Memorized
Pythia-1.4B Famous Quotes 1.17 0.47
Pythia-1.4B Paraphrased Quotes 0.68 0.11
@@ E.3 Paraphrased Famous Quotes
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Section 4
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations"section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We make no theoretical statement or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 2, Appendix B
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code included in the supplementary material and linked to in the footnote on the title page.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 4
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Our methods are for estimating the memorization of data by single models. For the illustrative purposes of our work, our empirical findings are averaged over datasets but not randomized trials. For this reason, we do not have error bars to report but we do discuss this in the limitations section.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes"if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix F
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification:
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Section 5
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: License is included when the data is introduced in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No human Subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.